# MaDiS: Taming Masked Diffusion Language Models for Sign Language Generation

### v1

@@ 4.2 Ablation Studies
We conduct all ablation studies on CSL-Daily.
@@
Tri-Level Cross-Modal Pretraining. We pretrain the model across three spaces (token, latent, and physical) to enable it to learn rich sign representations at different levels. Results in Table 3 validate the effectiveness of each pretraining level. First, at the token level, masking both sign and text tokens yields better performance than masking sign tokens alone, demonstrating the benefit of cross-modal training. This observation is consistent with previous findings on alignment-based pretraining in sign language translation [34, 92, 40]. Second, predicting latent features proves effective in reducing DTW errors and enhancing sentence-level semantic alignment. Finally, incorporating physical-space training by reconstructing sign motions leads to the best overall performance.
@@
6.61 9.05 3.87 3.11 32.81
$t=0.75$ 6.33 8.73 4.15 3.52 34.09
$t=0.5$ 6.18 8.23 4.63 3.94 36.73
$t=0.75/0.5$ 5.99 7.96 4.87 4.08 38.61
Table 4: Ablation study for unmasking with temporal checkpoints.
Unmasking with Temporal Checkpoints (UTC). To accelerate model convergence, we propose a novel unmasking strategy with temporal checkpoints, which substantially reduces the complexity of unmasking orders. As shown in Table 4, inserting checkpoints at either $t=0.75$ (corresponding to a $T/4$ temporal resolution) or $t=0.5$ ($T/2$ temporal resolution) yields superior performance compared to the no-checkpoint baseline. The best results are achieved when both checkpoints are applied. Furthermore, to examine whether the proposed UTC effectively accelerates training convergence, we visualize the training loss ($\mathcal{L}_{tok}$) and validation performance in Figure 6. The curves clearly demonstrate that incorporating checkpoints leads to faster convergence, which is reflected by lower training loss and higher validation performance at the same epoch.
Mixture-of-Parts (MoP) Embedding Layer. Our MoP sign embedding layer is built upon a dense mixture-of-experts (MoE) mechanism [10], in which all experts (body parts) are activated but controlled by a learnable gate. As shown in Table 5, we compare this design with a sparse MoE variant that fuses only the body parts corresponding to the top-1 or top-2 activation values. We observe that this sparse configuration results in inferior performance compared to the default dense setting. In Figure 7, we further visualize the learned gating weights for four example signs. For single-handed signs (left two examples), our MoP layer assigns higher weights to the active hand, indicating that it effectively captures hand dominance. In contrast, for two-handed signs, the layer learns comparable weights for both hands, reflecting balanced contributions during generation.
SiCLIP. To better evaluate the semantic alignment between texts and generated signs, we follow established practices in related fields [69, 42] and train a CLIP model over ground-truth sign motions and texts, referred to as SiCLIP. We assess its performance using the sign-to-text retrieval task. As shown in Table 6, simply using a contrastive loss [33] results in low retrieval performance. Introducing decoupled sign encoders [96] yields noticeable improvements. Most of the gains come from the fine-grained contrastive loss [16], which implicitly learns precise word-level alignments. The final retrieval performance is on par with that of a leading motion retrieval model [50], validating the effectiveness of SiCLIP as a reliable evaluator for future SLG research.
[…]
Body Hands Body Hands R@1
Simple Avg [96] 6.35 8.50 4.06 3.42 35.95
@@
Table 5: Ablation study for the MoP embedding layer.
@@
Baseline (CLIP) [33] 41.07 81.97 58.41 88.01 9.27 28.47
@@ 5 Conclusion
We present MaDiS, a novel SLG approach built upon MDLMs that enable bidirectional context modeling and efficient generation. To effectively inject sign language knowledge into MDLMs, we design a tri-level cross-modal pretraining framework that jointly learns from token, latent, and 3D physical spaces. To improve training efficiency of MDLMs, we simplify unmasking orders by proposing a novel unmasking strategy with temporal checkpoints. Moreover, a mixture-of-parts sign embedding layer is introduced to enhance part-wise token integration through gated, codebook-conditioned fusion. Experiments on standard SLG benchmarks demonstrate that MaDiS achieves SOTA performance across multiple evaluation metrics, while reducing inference latency by nearly 30%.
Acknowledgements. S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1), EPSRC Project GNOMON (EP/X011364/1) and Turing AI Fellowship (EP/Z534699/1). R.A. Potamias and R. Zuo were supported by EPSRC Project GNOMON (EP/X011364/1). J. Deng was supported by the NVIDIA Academic Grant.
The authors acknowledge the use of resources provided by the Isambard-AI National AI Research Resource (AIRR). Isambard-AI [41] is operated by the University of Bristol and is funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation; and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023].

### v2

@@ 4.2 Ablation Studies
We conduct all ablation studies on CSL-Daily unless otherwise specified.
Usage of Bidirectional Contexts. To verify that the MDLM effectively models bidirectional contexts, we visualize attention maps from the penultimate Transformer layer at the final sampling step, which captures rich contextual interactions while being less influenced by the task-specific output projection of the final layer [clark2019does, tenney2019bert_red]. As shown in Fig.˜7c, we select the three attention heads with the highest future-attention ratios, defined as $\sum_{i,j>i}\mathbf{A}_{i,j}/\sum_{i,j}\mathbf{A}_{i,j}$, where $\mathbf{A}$ denotes the attention map, and visualize their patterns. The results show that the model allocates substantial attention to future tokens (upper triangular region), indicating effective bidirectional modeling. Quantitatively, the average future-attention ratio reaches 41% across all heads over the test set.
@@
Figure 8: Visualization of reconstructed signs from mask tokens.
Tri-Level Cross-Modal Pretraining. We pretrain the model across three spaces (token, latent, and physical) to fully exploit the benefits from multi-level sign representations. Results in Tab.˜2 validate the effectiveness of each pretraining level. First, at the token level, masking both sign and text tokens yields better performance than masking sign tokens alone, demonstrating the benefit of cross-modal training. This observation is consistent with previous findings on alignment-based pretraining in sign language translation [jiao2024visual, zhou2023gloss, low2025sage]. Second, predicting latent features proves effective in reducing DTW errors and enhancing sentence-level semantic alignment. Finally, incorporating physical-space training by reconstructing sign motions leads to the best overall performance. This design also enables the pretrained model to learn physically grounded representations, allowing it to accurately reconstruct sign motions from masked tokens (Fig.˜8). Quantitatively, when $t=0.75$ (i.e., masking 25% of tokens), the reconstructed signs achieve mean position errors of 24.79/5.64mm on body/hand joints, comparable to the performance of the sign tokenizer [zuo2025soke].
[…]
(distribution, noise level) Body Hands Body Hands R@1
N/A 6.61 9.05 3.87 3.11 32.81
Uniform, $t=0.75$ 6.33 8.73 4.15 3.52 34.09
Uniform, $t=0.5$ 6.18 8.23 4.63 3.94 36.73
Uniform, $t=0.75/0.5$ 5.99 7.96 4.87 4.08 38.61
Contiguous, $t=0.75/0.5$ 6.28 8.44 4.31 3.80 34.85
Table 3: Ablation study for unmasking with temporal checkpoints.
Unmasking with Temporal Checkpoints (UTC). To accelerate model convergence, we propose a novel unmasking strategy with temporal checkpoints, which substantially reduces the complexity of unmasking orders in a coarse-to-fine manner. As shown in Tab.˜3, inserting checkpoints composed of uniformly distributed tokens at either $t=0.75$ ($T/4$ temporal resolution) or $t=0.5$ ($T/2$ temporal resolution) yields superior performance over the no-checkpoint setting. We further construct a baseline that retains only the first $1/4$ or $1/2$ contiguous tokens at the pivot states. While this design also reduces the order complexity, it does not explicitly model the coarse-to-fine prior. Although this baseline outperforms the no-checkpoint setting, the best performance is achieved when both checkpoints are applied with uniformly distributed tokens, indicating the benefits of both the reduced complexity and the coarse-to-fine design. Finally, we visualize the training loss ($\mathcal{L}_{tok}$) and validation performance in Fig.˜9. The curves demonstrate that our UTC leads to faster convergence, which is reflected by lower training loss and higher validation performance at the same epoch.
@@
N/A Average [zuo2025soke] 6.35 8.50 4.06 3.42 35.95
Codebook Average 6.19 8.25 4.45 3.73 36.98
MLP 6.22 8.29 4.19 3.86 37.02
Transformer 6.17 8.13 4.33 3.62 36.49
@@
Table 4: Ablation study for the MoP embedding.
@@
Category Left Right Two
Weights .41/.47/.12 .47/.08/.45 .32/.31/.37
Table 5: Weight distribution.
Mixture-of-Parts (MoP) Embedding Layer. Our MoP sign embedding layer dynamically fuses token embeddings from different body parts via a learnable gate. As shown in Sec.˜4.2, MoP significantly outperforms the prior design that simply averages randomly initialized part-wise token embeddings [zuo2025soke]. When conditioned on well-trained codebooks, we evaluate three standard fusion strategies: 1) averaging, 2) an MLP applied to the concatenation of part-wise code embeddings, and 3) a two-layer Transformer equipped with a $|\mathrm{CLS}|$ token. Results show that all conditioned fusion strategies outperform the non-conditioned baseline [zuo2025soke], validating the effectiveness of codebook conditioning.
Our MoP is built upon a dense mixture-of-experts (MoE) formulation [cai2025moe_survey], where all experts (corresponding to body parts) are activated and weighted by the gating network. As reported in Sec.˜4.2, we compare this design against sparse MoE variants that fuse only the body parts associated with the top-1 or top-2 gate activations. These sparse configurations consistently yield inferior performance relative to the default dense setting.
To further analyze the learned gating behavior, we visualize MoP gate weights for four example signs in Fig.˜10. For single-handed signs (left two examples), the MoP layer assigns higher weights to the active hand, indicating that it effectively captures hand dominance. In contrast, for two-handed signs, the layer learns comparable weights for both hands, reflecting balanced contributions during generation. Quantitatively, we identify the signing hand using a simple rule based on wrist position relative to the spine-01 joint [smplx], and accordingly classify signs as single- or two-handed. Tab.˜5 reports the correlation between MoP gate weights (body/left hand/right hand) and hand dominance, indicating that the learned gates align with interpretable physical structure.
@@
Baseline (CLIP) [jiang2024signclip] 41.07 81.97 58.41 88.01 9.27 28.47
@@
SiCLIP. To better evaluate the semantic alignment between texts and generated signs, we follow established practices in related fields [tevet2022motionclip, meng2025rethinking] and train a CLIP model over ground-truth sign motions and texts, referred to as SiCLIP. We assess its performance using the sign-to-text retrieval task. As shown in Tab.˜6, simply using a contrastive loss [jiang2024signclip] results in low retrieval performance. Introducing decoupled sign encoders [zuo2025soke] yields noticeable improvements. Most of the gains come from the fine-grained contrastive loss [cheng2023cico], which implicitly learns precise word-level alignments. The final retrieval performance is on par with that of a leading motion retrieval model [petrovich2023tmr], validating the reliability of SiCLIP as an SLG evaluator.
@@ 5 Conclusion
We present MaDiS, a novel SLG approach built upon MDLMs that enable bidirectional context modeling and efficient generation. To effectively inject sign language knowledge into MDLMs, we design a tri-level cross-modal pretraining framework that jointly learns from token, latent, and 3D physical spaces. To improve training efficiency of MDLMs, we simplify unmasking orders by proposing a novel unmasking strategy with temporal checkpoints. Moreover, a mixture-of-parts sign embedding layer is introduced to enhance part-wise token integration through gated, codebook-conditioned fusion. Experiments on standard SLG benchmarks demonstrate that MaDiS achieves SOTA performance, while improving inference throughput by 40%.
Acknowledgements. S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1), EPSRC Project GNOMON (EP/X011364/1) and Turing AI Fellowship (EP/Z534699/1). R.A. Potamias and R. Zuo were supported by EPSRC Project GNOMON (EP/X011364/1). J. Deng was supported by the NVIDIA Academic Grant. We would also like to thank Yuecong Min for coordinating the user study, and the anonymous signers who participated in it.
The authors acknowledge the use of resources provided by the Isambard-AI National AI Research Resource (AIRR). Isambard-AI [mcintosh2024isambard] is operated by the University of Bristol and is funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation; and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023].